OpenStudy (anonymous):

Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (anonymous):

what is the relative speed?

OpenStudy (anonymous):

???

OpenStudy (jim766):

distance = r t

OpenStudy (kainui):

distance is equal to rate times time, right? or written, d=rt. So in this case we're looking at their total distance which is the sum of both of their distances. They both have the same amount of time travelled at any point in time, right, the only difference is their distance and rate. So to tell the difference between the two I have given one a capital and lower case letters for their distance and rates. d=rt D=Rt So we can add both equations together. d+D=rt+Rt Now we can factor out "t" from the right side d+D=t(r+R) now we can divide out to solve for t, the time it takes. (d+D)/(r+R)=t and remember, d+D=45, r=10 and R=5. So now all you have to do is plug in your values!

OpenStudy (anonymous):

in this problem do I have to dd so subtract?

OpenStudy (anonymous):

ohh

OpenStudy (jim766):

45 = rt + rt

OpenStudy (anonymous):

don't make this more complicated than it needs to be the relative speed is 15 so the time is 45/15 = 3

OpenStudy (jim766):

45 = 10t + 5t

OpenStudy (jim766):

45 = 15t

OpenStudy (anonymous):

so that it you divide 45/15?

OpenStudy (jim766):

right...right....right

OpenStudy (anonymous):

oh kool allright :))) thanks you am stuck with another math problem , I already did the formula but still have to do something else. I f anyone can help me?? I will post it

OpenStudy (jim766):

post it...